# Statistics for Differential Topological Properties between Data Sets with an Application to Reservoir Computers

### v1

@@ IX Numerical examples of diffeomorphic statistics
In Eq.(16) $\kappa$ will be varied to change the dynamics, especially the RC damping. These variations will change the stability of both systems allowing us to examine the quality of the RCs reconstructions of the unseen Lorenz variables ($y$ and $z$) as stability varies up to the "edge of chaos" and compare those qualities simultaneously with the change in the continuity and differentiability statistic.
@@
IX.1 Continuity Statistic Test
@@
Later we discuss the case of a system which will not have a continuous function from the drive to the response system (a periodically driven period-doubled system). And will we see how the trends in Fig. 3 would be reversed indicating that there is no continuous function from the drive to the response.
@@
As we mentioned earlier the hyperparameter space for RCs is large. The parameters can be indirectly tuned for certain drive systems by altering the RC and seeing how well the it reconstructs the trajectories (as well as other RC tasks). However, a more direct way to probe the parameters that does not depend on an additional layer of dynamical calculations is to calculate the continuity statistic of the mapping between drive and RC. We can tune the parameters to guarantee the best statistical homeomorphism.
@@
Both extremes of dissipation relate directly to our former explanation of how the drive’s attractor is reconstructed in the RC and explain why too little or too much RC dissipation leads to poor performance of the RC. It is not that a fractal structure cannot be continuous image of another manifold, sometimes it is. The Weirstrass function is one example. But the amount of data necessary to overcome the "rough" fractal structure increases dramatically, hence the poor RC performance.
We note, that the differentiable nature of the drive-RC system will cease to be true in this case (as we will show below). No matter how much data we have or how the time series are sampled, a fractal structure is not a locally smooth manifold.
@@
IX.2 Existence of continuous functions into the future and past
@@
With the examination of the continuity statistic between time series into the past or future, we note that we can also think of the continuity statistic as a postdiction (into the past) and a prediction (into the future) test or, more accurately, function from one time in a dynamical time series to another. This helps us understand why the two plots in Fig. 5 track each other. The continuity statistic (expressed as $\epsilon^{*}/\epsilon_{min}$) basically shows that close points on the attractor at time shift=0 (the present) will be widely separated at times into the future and past. The training will not be constrained since there will be a large divergence of nearby points into the past and the future leading to poor data fitting of the output weights.
IX.3 Diffeomorphism Statistic
@@
$\sqrt{\sum_{i,j}\big{(}D\phi^{-1}D\phi-{\bf 1}\big{)}_{ij}^{2}/Q},$ (18)
where $Q$ is the local manifold dimension (see section VI).
@@
IX.4 Attractor Comparison Statistics and RC basins of attraction
Dynamical systems can have multiple attractors for some set of parameters, which includes driven systems. This means that what attractor the RC can settle into in different dynamical patterns in separate regions of the state space will depend on the initial conditions of the entire system, the drive and the RC. This was shown to occur in RC systems in the paper by Flynn et al.Flynn, Tsachouridis, and Amann (2021). Interestingly, in our case since the RC starts from an equilibrium state, where all the nodes are at the origin of its state space (all RC variables, ${\bf r}_{i}=0$) the final state depends on the drive system’s initial state, even though the drive still ends up in the same state. This situation is where our Attractor Comparison Statistic (ACS) will be helpful.
We run into the problem of multiple attractors when we have a Lorenz drive and a slightly different polynomial node RC system that we are using to train on reproduction of the $y$ and $z$ components from an $x$-driven RC and then test that reconstruction using a drive started at a different initial condition. The RC node dynamics are given by the same ODE system in Eq.16 with the overall coefficient $\kappa=1.0$ and we vary the damping coefficient $p_{1}$ in the polymomial over the range $-7.0$ to $-0.5$.
@@ IX.4 Attractor Comparison Statistics and RC basins of attraction
If we examine the RC node dynamics in pairs at points below, near, and above the dramatic testing error range we see very similar attractor shapes (which are, of course, projections from a higher 100 dimensional state space). These are shown in Fig.8.
@@
The results of the ACS are shown in Fig. 9. The ACS is small (approximately $10^{-2}$ or smaller for $p_{1}$ values at $-2.0$ and below for train-(time-shifted) train and test-train systems. But for larger $p_{1}$ values ($>-2.0$) the ACS for the comparison of train to test increases to above $10^{0}$, i.e. 100 times the other ACS more negative $p_{1}$ values. Meanwhile the train-(shifted) train ACS values remain small. Given the radical change in the trajectory of two components of each system’s coordinates at $p_{1}=-2.0$, we suspect that near that parameter value the RC system is undergoing a bifurcation at that $p_{1}$ value. This implied that the RC attractors were different in this parameter range.
@@
In examining in detail the trajectories in detail in each RC coordinate. We eventually found two coordinates in which the training and testing systems differed for $p_{1}>-2.0$ from the systems below that value. All other coordinates were the same. These are shown in Fig. 10. In this parameter region the plots of the system look similar (both are thin and slanted the same way, although the testing attractor appears to be tilted at a smaller (acute) angle.
@@
X Conclusions
The continuity and differentiability statistics in both directions gives two pairs of statistics (one for $\mathcal{D}\rightarrow\mathcal{R}$ and another for $\mathcal{R}\rightarrow\mathcal{D}$) that can give evidence that we either have a diffeomorphism or not between the drive and RC. As we mention above, the statistics can be used in more general situations to test for the existence (or not) of functions between more generally (paired) data sets. In a sense the continuity statistic suggest a function that predicts one set of data points from knowledge of another data set of points.
We addressed the case here where we have access to all dynamical variables. However, if we have a subset of variables that we can record, we can use a hybrid approach using delays in the recorded variables to reconstruct the state space trajectories and attractors. And, if the dimension of the drive is not too large, we will need to only reconstruct in a subspace of lower dimension to test for a diffeomorphism. Note that this approach may not always work. Or example using the ACS shows that sometimes a small number of RC nodes may have different dynamics from the trained dynamics. Then, curiously, the Takens reconstruction may not expose those. This situation needs to be studied in more detail.
We also showed that checking the differentiability statistic in both directions can be done with the diffeomorphic test for a possible local linear map between two (paired) data sets. But the test for a one-way differentiable function from one data set to the other can also be easily done and used when there is only a continuous function in one direction (see our period doubled counter-example earlier in this paper).
To do this we would gather local points from the function domain and their range twins. In our period doubled example the domain would be the period-two attractor and the range the period-one drive. As before we choose an even number of neighbors around each fiducial pt. After SVD tests for local tangent spaces, we then would use half of the points to calculate a linear function (matrix) representing a local linear transformation. We would calculate the error in using that matrix to map the 2nd half of the local domain points to their twins. A small error would imply a smooth map exists.
@@ X Conclusions
We finally note that the statistics are very general, especially the continuity statistic. We assume almost nothing about the data sets or the geometric structures they may be have. We note these advantages in the following list.
@@
$\bullet$ Derivative or diffeomorphism tests are useful in determining if fitting data or time series to some functional basis will be worthwhile.

### v2

@@ X Numerical examples of diffeomorphic statistics
With the exception of the ACS, the results presented here are all averages over 20 to 30 percent of the systems’ attractors, as determined from the sizes of the local sets of points used to calculate the various statistics. The statistics were calculated with increasing numbers of points until the results leveled off. The ACS was calculated using sets that covered 70% or more of the systems’ attractors. The standard deviations of all statistics were also calculated, but all showed no large variations in the the statistics’ groups and, hence, are not presented here.
@@
In Eq.(16) $\kappa$ will be varied to change the dynamics, especially the RC damping. These variations will change the stability of the RC allowing us to examine the quality of the RCs reconstructions of the unseen Lorenz variables ($y$ and $z$) as stability varies up to the "edge of chaos" and compare those qualities simultaneously with the change in the continuity and differentiability statistic.
@@
X.1 Continuity Statistic Test
@@
As we mentioned earlier the hyperparameter space for RCs is large. The parameters can be indirectly tuned for certain drive systems by altering the RC and seeing how well it reconstructs the trajectories (as well as other RC tasks). However, a more direct way to probe the parameters that does not depend on an additional layer of dynamical calculations is to calculate the continuity statistic of the mapping between drive and RC. We can tune the parameters to guarantee the best statistical homeomorphism.
@@
Both extremes of dissipation relate directly to our former explanation of how the drive’s attractor is reconstructed in the RC and explain why too little or too much RC dissipation leads to poor performance of the RC. It is not that a fractal structure cannot be continuous image of another manifold, sometimes it is. The Weierstrass function is one example (see Supplement A). But the amount of data necessary to overcome the "rough" fractal structure increases dramatically, hence the poor RC performance.
Note that the differentiable nature of the drive-RC system will cease to be true in this case (as we will show below). No matter how much data we have or how the time series are sampled, a fractal structure is not a locally smooth manifold. This means the RC will not replicate the dimension of the drive or the drive’s Lyapunov exponents. We show an example of this in Supplement A.
Fig-TimeShift-Resv-to-Lor-cnt-fits/Fig-TimeShift-cnt-fit-R esv to-Lor-ppt
@@
X.2 Existence of continuous functions into the future and past
@@
With the examination of the continuity statistic between time series into the past or future, we can also think of the continuity statistic as a postdiction (into the past) and a prediction (into the future) test or, more accurately, a function from one time in a dynamical time series to another. This helps us understand why the two plots in Fig. 5 track each other. The good continuity into the future means we can successfully use the training to predict the values of the dynamical variables into the near future making it possible to use the RC as a autonomous dynamical system Lu, Hunt, and Ott (2018); Kong, Brewer, and Lai (2024).
The continuity statistic also shows that close points on the attractor at time shift=0 (the present) will be widely separated at times into the more distant future and past. The training will not be constrained since there will be a large divergence of nearby points into the past and the future leading to poor data fitting of the output weights.
X.3 Diffeomorphism Statistic
@@
$\sqrt{\sum_{i,j}\big{(}D\phi^{-1}D\phi-{\bf 1}\big{)}_{ij}^{2}/P},$ (18)
where $P$ is the local manifold dimension (see section VI).
@@
X.4 Attractor Comparison Statistics and RC basins of attraction
Dynamical systems can have multiple attractors for some set of parameters, which includes driven systems. This means that what attractor the RC can settle into in different dynamical patterns in separate regions of the state space will depend on the initial conditions of the entire system, the drive and the RC. This was shown to occur in RC systems in the paper by Flynn et al.Flynn, Tsachouridis, and Amann (2021). Interestingly, in our case since the RC starts from an equilibrium state, where all the nodes are at the origin of its state space (all RC variables, ${\bf r}_{i}=0$), but the final state depends on the drive system’s initial state, even though the drive still ends up in the same state. This situation is where our Attractor Comparison Statistic (ACS) will be helpful.
We run into this problem of multiple attractors when we have a Lorenz drive and a slightly different polynomial node RC system that we are using to train on reproduction of the $y$ and $z$ components from an $x$-driven RC and then test that reconstruction using a drive started at a different initial condition. The RC node dynamics are given by the same ODE system in Eq.16 with the overall coefficient $\kappa=1.0$ and we vary the damping coefficient $p_{1}$ in the polymomial over the range $-7.0$ to $-0.5$.
@@ X.4 Attractor Comparison Statistics and RC basins of attraction
If we examine some of the 100 RC nodes’ dynamics in pairs of components at points below, near, and above the dramatic testing error range we see very similar attractor shapes (which are, of course, projections from a higher 100 dimensional state space). These are shown in Fig.8.
@@
The results of the ACS are shown in Fig. 9. The ACS is small (approximately $10^{-2}$ or smaller for $p_{1}$ values at $-2.0$ and below for train-(time-shifted) train and test-train systems. But for larger $p_{1}$ values ($>-2.0$) the ACS for the comparison of train to test increases to above $10^{0}$, i.e. 100 times the other ACS at more negative $p_{1}$ values. Meanwhile the train-(shifted) train ACS values remain small. Given the radical change in the trajectory of the components of each system’s coordinates at $p_{1}=-2.0$, we suspect that near that parameter value the RC system is undergoing a bifurcation at that $p_{1}$ value. This implied that the RC attractors were different in this parameter range.
@@
In examining in detail the trajectories in detail in each RC coordinate. We eventually found two coordinates in which the training and testing systems differed for $p_{1}>-2.0$ from the systems below that value. All other coordinates were the same. These two are shown in Fig. 10. In this parameter region the plots of the system look similar (both are thin and slanted the same way, although the testing attractor appears to be tilted at a smaller (acute) angle.
@@
XI Conclusions
The continuity and differentiability statistics in both directions gives two pairs of statistics (one for $\mathcal{D}\rightarrow\mathcal{R}$ and another for $\mathcal{R}\rightarrow\mathcal{D}$) that can give evidence that we either have a diffeomorphism or not between the drive and RC. As we mention above, the statistics can be used in more general situations to test for the existence (or not) of functions between more generally (paired) data sets. In a sense the continuity statistic suggests a function exists that predicts one set of data points from knowledge of another data set of points.
We addressed the case here where we have access to all dynamical variables. However, if we have a subset of variables that we can record, we can use a hybrid approach using delays in the recorded variables to reconstruct the state space trajectories and attractors. And, if the dimension of the drive is not too large, we will need to only reconstruct in a subspace of lower dimension to test for a diffeomorphism. Note that this approach may not always work. Our example using the ACS shows that sometimes a small number of RC nodes may have different dynamics from the trained dynamics. Then, curiously, the Takens reconstruction may not expose those. This situation needs to be studied in more detail.
We also showed that checking the differentiability statistic in both directions can be done with the diffeomorphic test for a possible local linear map between two (paired) data sets. But the test for a one-way differentiable function from one data set to the other can also be easily done, too and used when there is only a continuous function in one direction (see our period doubled counter-example earlier in this paper).
@@ XI Conclusions
We finally note that the statistics we present here are very general, especially the continuity statistic. We assume almost nothing about the data sets or the geometric structures they possibly have. We list these advantages as follows,
@@
$\bullet$ Derivative or diffeomorphism tests are useful in determining if fitting data or time series to some smooth functional basis will be worthwhile.
Supplement A: Continuity and Differentiability of the Weierstrass (Fractal) Function.
The Weierstrass function is a formula for approximating a real continuous, but non-differentiable function on the real line with a type of Fourier series as followsMandelbrot (1983); Wikipedia contributors (2024):
$w(x)={\sum_{n=0}^{\infty}a^{n}\cos(b^{n}\pi x)},$ (19)
where $a=0.7$ and $b=10.0$ and $x\in[0,1]$.
Since we will be generating computed numbers we cut off the sum at a number $N$ for two cases: (1) $N=2$ to generate numerically differentiable function variables and (2) $N=20$ for numerically non-differentiable function variables.
Figure 11: Smooth version of the "Weierstrass function," $N=2$ version.
Figure 12: Fractal version of the Weierstrass function, $N=20$ version with small portion of $w(x)$ expanded to show more structure in small regions.
The plot of case (1) "Weierstrass" function is shown in Fig. 11 and it is easy to see that it is smooth, i.e. non-fractal.
For case (2) the large $N$ value guarantees that the fluctuations in the trigonometric function are below the computer double precision ($\sim 10^{-16}$) and the minimum spacing between points in the numerics below is roughly $10^{-4}$. In this case the numerical calculations will "see" the function values as continuous, but non-differentiable. Fig. 12 shows this in some detail.
In Fig. 13 we plot the continuity statistics for the $N=2$ and $N=20$ cases. For the N=2 case the $\epsilon^{*}/\epsilon_{min}$ ratio stays constant as we add more points to the time series, i.e. increasing the density of points that are sampled for the statistic. The actual values of $\epsilon^{*}$ and $\epsilon_{min}$ decrease with increasing number of data points. These results for the continuity statistic provide evidence that the $N=2$ function is continuous.
In the same figure we plot the continuity statistic for the fractal ($N=20$) Weierstrass case. Although the $\epsilon^{*}$ statistic does not fall as fast as $\epsilon_{min}$, it does decrease with increasing density of domain points. This suggests a continuous function, but one with "fine" structure.
Figure 13: Continuity statistic for "smooth" and "fractal" versions of the Weierstrass function. The continuity statistic $\epsilon^{*}$ decreases with increase density of points, but not as quickly as the $\epsilon_{min}$ statistic. The differing rates of decrease are related to the fractal structure of the function.
Finally, the derivative statistic for the two $N$ cases show that the $N=2$ case behaves like a differentiable function, but the $N=20$ case has large errors (near 50%), which do not decrease at all even was we add an order of magnitude more points to the same interval.
Figure 14: The Derivative statistic for smooth and fractal Weierstrass functions. The $N=2$ smooth Weierstrass $D\phi$ error asymptotically becomes linear and small as one would expect when the nearest neighbors of the function points are well approximated with the slope. But the $N=20$ Weierstrass $D\phi$ error starts out at an error of 57% (a large error) and never deviates from values in that range, implying that the true Weierstrass function is not differentiable.